Elvira drives into the city to see a play. Due to traffic, she can only travel at arate of 15 mph. On her ride home, she is able to average 35 mph. Her total travel is 2 hours. How long does it take her to get home?
Answer by Greenfinch(383)

Time in =distance/15
Time back = distance/35
So total time = 2 = d/15 + d/35 = 7d/105 + 3d/105 = 10d/105
So 210 = 10 d or d = 21 miles
Now we know the distance, the times can be worked out
Time out = 21/15 = 1.4 hours
Time in = 21/35 = 0.6 hours
Total 2 hours, time to get back, 0.6 hours